OpenStudy (anonymous):

Determine which ordered pair in the form (a, b) is a solution of the given equation. 7a – 5b = 28 (Points : 1) (–3, –2) (–2, –3) (0, 4) (4, 0)

OpenStudy (anonymous):

\[7a – 5b = 28\] if \(a=4,b=0\) you have \[7\times 4-5\times 0=28\] so that one works maybe some others as well

OpenStudy (anonymous):

nah, i think that is the only one

OpenStudy (confusionist):

Yeah, I just checked; That's the only one^

OpenStudy (anonymous):

Thank you (: